OpenStudy (crashonce):

THe interior angles of an n-sided convex polygon are in AP. If the common difference is 5 and the largest angle is 160, then n equals? @ganeshie8

ganeshie8 (ganeshie8):

lets try and setup two equations in "n" and "a"

ganeshie8 (ganeshie8):

largest angle is 160 :\[a+(n-1)5 = 160\tag{1}\]

ganeshie8 (ganeshie8):

interior angles add up to 180(n-2) : \[\dfrac{n}{2}\left[2a+(n-1)5\right] = 180(n-2)\tag{2}\]

ganeshie8 (ganeshie8):

two equations and two unknowns, you can solve them

OpenStudy (crashonce):

one sec

ganeshie8 (ganeshie8):

its a bit trickier than the previous problem as you can see the second equation is quadratic, so you will need to do a bit more work

OpenStudy (crashonce):

cant get it

OpenStudy (crashonce):

show me

ganeshie8 (ganeshie8):

try to substitute the value of "a" from first equation into second equation

OpenStudy (crashonce):

i tried but it doesnt seem to work

OpenStudy (crashonce):

wait

ganeshie8 (ganeshie8):

ok waiting

OpenStudy (crashonce):

is it 16

OpenStudy (crashonce):

@ganeshie8

ganeshie8 (ganeshie8):

doesn't look correct

ganeshie8 (ganeshie8):

isolating "a" from first eqn gives you \[a = 160-5(n-1)\] substitute that in second eqn

ganeshie8 (ganeshie8):

\[\dfrac{n}{2}\left[2(160-5(n-1))+(n-1)5\right] = 180(n-2)\] simplify a bit and you will see a nice quadratic in "n"

OpenStudy (crashonce):

what is it

OpenStudy (crashonce):

i cant seem to get it

ganeshie8 (ganeshie8):

have you understood everythign so far ?

OpenStudy (crashonce):

yes

ganeshie8 (ganeshie8):

Okay do you know to use distributive property and combine like terms ?

OpenStudy (crashonce):

yes

OpenStudy (crashonce):

i just cant gettit for some strange reason

ganeshie8 (ganeshie8):

then whats stopping you from simplifying the above mess ?

ganeshie8 (ganeshie8):

\[\dfrac{n}{2}\left[2(160-5(n-1))+(n-1)5\right] = 180(n-2)\] start by distrubuting \[\dfrac{n}{2}\left[2\times 160-2\times 5(n-1))+5(n-1)\right] = 180\times n-180\times 2\]

OpenStudy (crashonce):

i know i didnt get a quadratic but got n=18 which is wrong

ganeshie8 (ganeshie8):

\[\dfrac{n}{2}\left[2\times 160-5(n-1)\right] = 180n-360\]

ganeshie8 (ganeshie8):

\[\dfrac{n}{2}\left[325-5n\right] = 180n-360\]

ganeshie8 (ganeshie8):

multiply 2 both sides \[n\left[325-5n\right] = 360n-720\]

OpenStudy (crashonce):

the roots are -16 and 9

OpenStudy (crashonce):

so its 9 oooops forgot to redo the minus

ganeshie8 (ganeshie8):

Yep! discard -16 as the number of sides cannot be negative

ganeshie8 (ganeshie8):

The key part is setting up the equations correctly using the properties of AP and triangles. Once you have the equations, you can solve them using any of your favorite methods